# COADVISE: Covariate Adjustment with Variable Selection in Randomized Controlled Trials

### v4

@@ X The covariate vector/matrix/data frame $\checkmark$
Figure S.2.10: Simulation results under binary outcome, linear $\delta(\mathbf{X})$ and $N=100$. In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels. In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
@@
Figure S.2.11: Simulation results under binary outcome, linear $\delta(\mathbf{X})$ and $N=200$. In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels. In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
@@
Figure S.2.12: Simulation results under binary outcome, linear $\delta(\mathbf{X})$ and $N=500$. In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels. In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
@@
Figure S.2.13: Simulation results under binary outcome, nonlinear $\delta(\mathbf{X})$ and $N=40$. In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels. In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
@@
Figure S.2.14: Simulation results under binary outcome, nonlinear $\delta(\mathbf{X})$ and $N=100$. In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels. In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
@@
Figure S.2.15: Simulation results under binary outcome, nonlinear $\delta(\mathbf{X})$ and $N=200$. In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels. In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
@@
Figure S.2.16: Simulation results under binary outcome, nonlinear $\delta(\mathbf{X})$ and $N=500$. In the CPplots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels. In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.

### v5

@@ X The covariate vector/matrix/data frame $\checkmark$
Figure S.2.10: Simulation results under binary outcome, linear $\delta(\mathbf{X})$ and $N=100$. In the CP% plots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels. In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
@@
Figure S.2.11: Simulation results under binary outcome, linear $\delta(\mathbf{X})$ and $N=200$. In the CP% plots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels. In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
@@
Figure S.2.12: Simulation results under binary outcome, linear $\delta(\mathbf{X})$ and $N=500$. In the CP% plots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels. In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
@@
Figure S.2.13: Simulation results under binary outcome, nonlinear $\delta(\mathbf{X})$ and $N=40$. In the CP% plots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels. In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
@@
Figure S.2.14: Simulation results under binary outcome, nonlinear $\delta(\mathbf{X})$ and $N=100$. In the CP% plots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels. In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
@@
Figure S.2.15: Simulation results under binary outcome, nonlinear $\delta(\mathbf{X})$ and $N=200$. In the CP% plots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels. In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.
@@
Figure S.2.16: Simulation results under binary outcome, nonlinear $\delta(\mathbf{X})$ and $N=500$. In the CP% plots, the red dashed line indicates 95% coverage level, and the two gray dashed lines indicate 93% and 97% coverage levels. In the power plots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power.